OpenStudy (anonymous):

if cosx=u/sq(u^2+1), find expression for sin2x in terms of u

OpenStudy (anonymous):

\[\cos x=\frac{ u }{ \sqrt{u^2+1} }\]

OpenStudy (anonymous):

Sin2x = 2 sinx cosx sin^2 x + cos^2 x =1

OpenStudy (anonymous):

u have cos x find sinx first.....using sin^2 x + cos^2 x =1 and subs the value in Sin2x = 2 sinx cosx

OpenStudy (anonymous):

can you please explain further